### v2

Abstract.
The $n$-dimensional hypercube $Q_{n}$ is a graph with vertex set $\{0,1\}^{n}$ and there is an edge between two vertices if they differ in exactly one coordinate. For any graph $H$, define $\textup{ex}(Q_{n},H)$ to be the maximum number of edges of a subgraph of $Q_{n}$ without a copy of $H$. In this short note, we prove that for any $n\in\mathbb{N}$
$\textup{ex}(Q_{n},C_{10})>0.024\cdot e(Q_{n}).$
Our construction is strongly inspired by the recent breakthrough work of Ellis, Ivan, and Leader k[12], who showed that “daisy” hypergraphs have positive Turán density with an extremely clever and simple linear-algebraic argument.
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior, Brasil (CAPES)
1. Introduction
For each $n\in\mathbb{N}$, define the $n$-dimensional hypercube $Q_{n}$ to be the graph with vertex set $\{0,1\}^{n}$, and with an edge between two vertices if they differ in exactly one coordinate. Erdős [13] initiated the study of $\textup{ex}(Q_{n},H)$, the maximum number of edges in an $H$-free subgraph of $Q_{n}$, in the special case where $H$ is an even cycle. We say that $H$ has positive Turán density in the hypercube if there is some constant $\alpha>0$ such that for every $n\in\mathbb{N}$
$\textup{ex}(Q_{n},H)\geqslant\alpha\cdot e(Q_{n}),$
where $e(Q_{n})$ denotes the number of edges of $Q_{n}$.
It will be useful to identify in the usual way the vertices of $Q_{n}$ with the subsets of $[n]:=\{1,\ldots,n\}$. For any $r\leqslant n$, we define $L_{r}(n)$, the $r$-th (edge) layer of $Q_{n}$, to be the subgraph of $Q_{n}$ formed by the edges between $\binom{[n]}{r-1}$ and $\binom{[n]}{r}$, where
$\binom{[n]}{r}:=\big{\{}S\subset[n]:|S|=r\big{\}}.$
It is easy to see that $C_{4}\not\subset L_{r}(n)$, thus taking every second layer gives that
$\textup{ex}(Q_{n},C_{4})\geqslant\frac{1}{2}\cdot e(Q_{n}).$
Erdős [14] offered \$$\$100$ for an answer to whether this bound is optimal, and it remains open until the present date. The best known upper bound, due to Baber [4], is approximately $0.60318\cdot e(Q_{n})$ (see [9, 17, 5] for previous upper bounds).
Erdős [13] suggested that longer even cycles might not have positive Turán density, but this was shown to be false for $C_{6}$ by Chung [9] and Brouwer et al. [7]. On the other hand, it was proved by Chung [9] for every even $t\geqslant 4$, and by Füredi and Özkahya [15, 16] for every odd $t\geqslant 7$, that
$\textup{ex}(Q_{n},C_{2t})=o(e(Q_{n})),$
and a unified proof for all of the above cases was given by Conlon [11] (the best known upper bounds for $t$ large are due to Chung [9] when $t$ is even, and Tomon [18, Theorem 4.1] when $t$ is odd). Thus, the only case for which the problem remained open was $C_{10}$. The main result of this paper completes the picture, by showing that $C_{10}$ has positive Turán density in the hypercube.
Theorem 1.1.
For all $n\in\mathbb{N}$
$\textup{ex}(Q_{n},C_{10})>0.024\cdot e(Q_{n}).$
In fact, our proof gives the constant $c/12$, where
$c=\prod_{k=1}^{\infty}\bigg{(}1-\frac{1}{2^{k}}\bigg{)}>0.288.$
We also remark that the graph we construct to prove Theorem 1.1 is free of $C_{4}$, $C_{6}$ and $C_{10}$ simultaneously. The previous best known lower bound in this case was
$\textup{ex}(Q_{n},C_{10})\geqslant\frac{e(Q_{n})}{(\log n)^{\alpha}}$
for some constant $\alpha>0$, recently shown by Axenovich et al. [3, Theorem 5].
The Ramsey problem for even cycles in the hypercube has also attracted a great deal of attention over the years. In particular, Chung [9] and Brouwer, Dejter, and Thomassen [7] found $4$-colourings of the hypercube without a monochromatic $C_{6}$, and Conder [10] found a $3$-colouring with the same property, implying the best known lower bound for this case of $\textup{ex}(n,C_{6})\geqslant\frac{1}{3}e(Q_{n})$. However, it was shown by Alon et al. [1] that for any $k$-colouring of the edges of $Q_{n}$ there is a monochromatic copy of $C_{10}$. Our result therefore gives, to the best of our knowledge, the first example of a Ramsey graph which has positive Turán density in the hypercube. The best upper bound we are aware of is $\textup{ex}(Q_{n},C_{10})\leqslant\frac{1}{\sqrt{2}}\cdot e(Q_{n})$ by Axenovich and Martin [2, Theorem 3.3].
Our proof of Theorem 1.1 is heavily inspired by a recent breakthrough work of Ellis, Ivan, and Leader [12], who gave an extremely clever and simple linear-algebraic construction which shows that “daisies” have positive Turán density in hypergraphs, disproving a conjecture of Bollobás et al. [6] and Bukh [8]. The argument we use is essentially a slight modification of their approach.
We will need another important idea, first observed by Alon, Radoičić, Sudakov, and Vondrák [1] and Axenovich and Martin [2]: it is possible to find a $4$-colouring of $E(Q_{n})$ without monochromatic chordless $C_{10}$. More precisely, define $\textup{ex}^{*}(Q_{n},C_{6}^{-})$ to be the maximum number of edges of a subgraph of $Q_{n}$ such that there is no embedding of $C_{6}$ minus one edge that can be extended to a $C_{6}$ in $Q_{n}$. We will make use of the following result of Axenovich, Martin, and Winter [3], which is based on the $3$-colouring given by Conder [10].
Theorem 1.2 (Lemma 19 in [3]).
$\textup{ex}(Q_{n},C_{10})\geqslant\frac{1}{3}\cdot\textup{ex}^{*}(Q_{n},C_{6}^{-}).$
Thus, the core of the proof will be to show the following theorem.
Theorem 1.3.
For any $r,n\in\mathbb{N}$ with $r\leqslant n$, there exists a $C_{6}$-free induced subgraph $G_{r}$ of $L_{r}(n)$ with
$e(G_{r})>\frac{c}{2}\cdot e(L_{r}(n)).$
Proof of Theorem 1.1 from Theorem 1.3.
Take $G$ to be the union of the graphs $G_{r}$ given by Theorem 1.3 for every odd $r\leqslant n$, so that $G_{r}$ is disconnected from $G_{r^{\prime}}$ for every $r\neq r^{\prime}$.
Observe that every induced subgraph of a layer containing a $C_{6}^{-}$ also contains a $C_{6}$; thus, since $G_{r}$ is a $C_{6}$-free induced subgraph of $L_{r}(n)$, it is also free of $C_{6}^{-}$. This implies
$\textup{ex}^{*}(Q_{n},C_{6}^{-})\geqslant e(G)>\frac{c}{2}\sum_{r=1}^{\left\lceil n/2\right\rceil}e\big{(}L_{2r-1}(n)\big{)}=\frac{c}{4}\cdot e(Q_{n}),$
which, by Theorem 1.2, concludes the proof. ∎
2. Proof of Theorem 1.3
Similarly to the argument used in [12], we consider the vector space $\mathbb{F}_{2}^{r}$. Also we fix a nonzero vector $v_{0}\in\mathbb{F}_{2}^{r}$. For each $i\in[n]$, pick a uniform random vector $v_{i}\in\mathbb{F}_{2}^{r}\setminus\{0\}$. Now each vertex of our hypercube layer $L_{r}(n)$ corresponds to a certain collection of vectors in $\mathbb{F}_{2}^{r}$ (of size $r$ or $r-1$). Define the sets
$B_{r}=\bigg{\{}S\in{[n]\choose r}:\big{\{}v_{i}:i\in S\big{\}}\text{ form a basis of }\mathbb{F}_{2}^{r}\bigg{\}},$
$B_{r-1}=\bigg{\{}S\in{[n]\choose{r-1}}:\big{\{}v_{i}:i\in S\big{\}}\cup\{v_{0}\}\text{ form a basis of }\mathbb{F}_{2}^{r}\bigg{\}},$
and consider the induced subgraph $G_{r}$ on $B_{r-1}\cup B_{r}$.
We will show that this (random) graph $G_{r}$ is $C_{6}$-free (deterministically) and has a large expected number of edges, and therefore has the desired properties with positive probability.
Claim 2.1.
$G_{r}$ is $C_{6}$-free.
Proof.
Suppose there is a copy of $C_{6}$ in $G_{r}$. There are exactly three different coordinates flipped by its edges, so it must form the middle layer of some $3$-dimensional subcube of $Q_{n}$. More precisely, by permuting the coordinates, we may assume that this copy of $C_{6}$ is exactly of the form $L_{r}(n)\big{[}A_{1}\cup A_{2}\big{]}$, where
$A_{1}=\big{\{}\{i\}\cup I:1\leqslant i\leqslant 3\big{\}}\qquad\text{and}\qquad A_{2}=\big{\{}\{i,j\}\cup I:1\leqslant i<j\leqslant 3\big{\}}$
for some $I\subset[n]\setminus\{1,2,3\}$ with $|I|=r-2$. Also let $U=\{v_{i}:i\in I\}$ be the collection of vectors corresponding to the coordinates in $I$.
Observe that, since $\{1,2\}\cup I\in A_{2}\subset B_{r}$, the collection of vectors $\{v_{1},v_{2}\}\cup U$ is a basis of $\mathbb{F}_{2}^{r}$, so, in particular, the vectors in $U$ are linearly independent. Therefore, after taking the quotient of $\mathbb{F}_{2}^{r}$ by $\langle U\rangle$ we obtain a vector space $V$ isomorphic to $\mathbb{F}_{2}^{2}$. Define $x_{0},x_{1},x_{2},x_{3}\in V$ to be the images of $v_{0},v_{1},v_{2},v_{3}$, respectively, under the quotient map. Since $A_{2}\subset B_{r}$ we have that
$\{v_{i},v_{j}\}\cup U\text{ form a basis of $\mathbb{F}_{2}^{r}$ for any $1\leqslant i<j\leqslant 3$,}$
i.e.,
[…]
Similarly, from $A_{1}\subset B_{r-1}$, by the same argument as above, we obtain
$\{x_{0},x_{i}\}\text{ form a basis of $V$ for any $1\leqslant i\leqslant 3$.}$
In particular, this implies that $x_{i}\neq 0$ for each $i\in\{0,1,2,3\}$ and that $x_{i}\neq x_{j}$ for any $i\neq j$. But $|V\setminus\{0\}|=3$, which yields a contradiction. ∎
Claim 2.2.
$\mathbb{E}\big{[}e(G_{r})\big{]}>\dfrac{c}{2}\cdot e\big{(}L_{r}(n)\big{)}$.
Proof.
Consider an edge of $L_{r}(n)$ connecting two sets $x=\{j_{1},\ldots,j_{r-1}\}$ and $y=\{j_{1},\ldots,j_{r}\}$. Define the vector spaces
$V_{k}=\langle v_{0},v_{j_{1}},\ldots,v_{j_{k}}\rangle\quad\text{ for $0\leqslant k\leqslant r-1$, and}\quad V_{r}=\langle v_{j_{1}},\ldots,v_{j_{r}}\rangle.$
Note that
$\mathbb{P}\big{(}xy\in G_{r}\big{)}=\mathbb{P}\big{(}x\in B_{r-1}\text{ and }y\in B_{r}\big{)}=\mathbb{P}\big{(}\dim V_{r-1}=\dim V_{r}=r\big{)}.$ (1)
Observe that, if the vectors $v_{0},v_{j_{1}},\ldots,v_{j_{k-1}}$ (for $11\leqslant k\leqslant r-$ 1) are fixed and linearly independent, then there are exactly $2^{r}-2^{k}$ admissible choices for the next vector $v_{j_{k}}$. Similarly, if the vectors $v_{j_{1}},\ldots,v_{j_{r-1}}$ are fixed and linearly independent, then there are $2^{r}-2^{r-1}$ admissible choices for the last vector $v_{j_{r}}$. Therefore, by (1), we have
$\mathbb{P}\big{(}xy\in G_{r}\big{)}$ $=\Bigg{(}\prod_{k=1}^{r-1}\mathbb{P}\big{(}v_{j_{k}}\notin V_{k-1}\;|\;\dim V_{k-1}=k\big{)}\Bigg{)}\cdot\mathbb{P}\big{(}v_{j_{r}}\notin\langle v_{j_{1}},\ldots,v_{j_{r-1}}\rangle\;|\;\dim V_{r-1}=r\big{)}$
$=\Bigg{(}\prod_{k=1}^{r-1}\frac{2^{r}-2^{k}}{2^{r}-1}\Bigg{)}\cdot\frac{2^{r}-2^{r-1}}{2^{r}-1}>\frac{1}{2}\prod_{k=1}^{\infty}\bigg{(}1-\frac{1}{2^{k}}\bigg{)}=\frac{c}{2}.$
The claim now follows from linearity of expectation. ∎
As observed above, it follows from Claims 2.1 and 2.2 that there exists a choice of the vectors $v_{1},\ldots,v_{n}$ such that
$e(G_{r})>\frac{c}{2}\cdot e\big{(}L_{r}(n)\big{)}$
and $G_{r}$ is $C_{6}$-free, as required.                                                                                                                                     ∎