### v3

Abstract.
The $n$-dimensional hypercube $Q_{n}$ is a graph with vertex set $\{0,1\}^{n}$ such that there is an edge between two vertices if and only if they differ in exactly one coordinate. For any graph $H$, define $\textup{ex}(Q_{n},H)$ to be the maximum number of edges of a subgraph of $Q_{n}$ without a copy of $H$. In this short note, we prove that for any $n\in\mathbb{N}$
$\textup{ex}(Q_{n},C_{10})>0.024\cdot e(Q_{n}),$
where $e(Q_{n})$ is the number of edges of $Q_{n}$. Our construction is strongly inspired by the recent breakthrough work of Ellis, Ivan, and Leader, who showed that “daisy” hypergraphs have positive Turán density with an extremely clever and simple linear-algebraic argument.
Key words and phrases:
extremal graph theory, hypercube, even cycles
1. Introduction
For each $n\in\mathbb{N}$, define the $n$-dimensional hypercube $Q_{n}$ to be the graph with vertex set $\{0,1\}^{n}$ such that there is an edge between two vertices if and only if they differ in exactly one coordinate. Erdős [13] initiated the study of $\textup{ex}(Q_{n},H)$, the maximum number of edges in an $H$-free subgraph of $Q_{n}$, in the special case when $H$ is an even cycle. We say that $H$ has positive Turán density in the hypercube if there is some constant $\alpha>0$ such that for every $n\in\mathbb{N}$
$\textup{ex}(Q_{n},H)\geqslant\alpha\cdot e(Q_{n}),$
where $e(Q_{n})$ denotes the number of edges of $Q_{n}$.
We identify the vertices of $Q_{n}$ with the subsets of $[n]:=\{1,\ldots,n\}$ in the usual way. For any $r\leqslant n$, we define $L_{r}(n)$, the $r$-th (edge) layer of $Q_{n}$, to be the subgraph of $Q_{n}$ formed by the edges between $\binom{[n]}{r-1}$ and $\binom{[n]}{r}$, where
$\binom{[n]}{r}:=\big{\{}S\subset[n]:|S|=r\big{\}}.$
It is easy to see that $C_{4}\not\subset L_{r}(n)$, thus taking every second layer gives that
$\textup{ex}(Q_{n},C_{4})\geqslant\frac{1}{2}\cdot e(Q_{n}).$
Erdős [14] offered \$$\$100$ for an answer to whether this bound is optimal, and it remains open until the present date. The best known upper bound, due to Baber [4], is approximately $0.60318\cdot e(Q_{n})$ (see [9, 17, 5] for previous upper bounds).
Erdős [13] suggested that longer even cycles might not have positive Turán density, but this was shown to be false for $C_{6}$ by Chung [9] and Brouwer et al. [7]. On the other hand, it was proved by Chung [9] for every even $t\geqslant 4$, and by Füredi and Özkahya [15, 16] for every odd $t\geqslant 7$, that
$\textup{ex}(Q_{n},C_{2t})=o(e(Q_{n})),$
and a unified proof for all of the above cases was given by Conlon [11] (the best known upper bounds for $t$ large are due to Chung [9] when $t$ is even, and Tomon [18, Theorem 4.1] when $t$ is odd). Thus, the only case for which the problem remained open was $C_{10}$. The main result of this paper completes the picture, by showing that $C_{10}$ has positive Turán density in the hypercube.
Theorem 1.1.
For all $n\in\mathbb{N}$
$\textup{ex}(Q_{n},C_{10})>\frac{c}{12}\cdot e(Q_{n}),$
where
$c=\prod_{k=1}^{\infty}\bigg{(}1-\frac{1}{2^{k}}\bigg{)}>0.288.$
We remark that the graph we construct to prove Theorem 1.1 is free of $C_{4}$, $C_{6}$ and $C_{10}$ simultaneously. The previous best known lower bound in this case was
$\textup{ex}(Q_{n},C_{10})\geqslant\frac{e(Q_{n})}{(\log n)^{\alpha}}$
for some constant $\alpha>0$, recently shown by Axenovich et al. [3, Theorem 4].
The Ramsey problem for even cycles in the hypercube has also attracted a great deal of attention over the years. In particular, Chung [9] and Brouwer, Dejter, and Thomassen [7] found $4$-colourings of the hypercube without a monochromatic $C_{6}$, and Conder [10] found a $3$-colouring with the same property, implying the best known lower bound for this case of $\textup{ex}(n,C_{6})\geqslant\frac{1}{3}e(Q_{n})$. However, it was shown by Alon et al. [1] that for any $k$-colouring of the edges of $Q_{n}$ there is a monochromatic copy of $C_{10}$. Our result therefore gives, to the best of our knowledge, the first example of a Ramsey graph which has positive Turán density in the hypercube. The best upper bound we are aware of is $\textup{ex}(Q_{n},C_{10})\leqslant\frac{1}{\sqrt{2}}\cdot e(Q_{n})$ by Axenovich and Martin [2, Theorem 3.3].
Our proof of Theorem 1.1 is heavily inspired by a recent breakthrough work of Ellis, Ivan, and Leader [12], who gave an extremely clever and simple linear-algebraic construction which shows that “daisies” have positive Turán density in hypergraphs, disproving a conjecture of Bollobás et al. [6] and Bukh [8]. The argument we use is essentially a slight modification of their approach.
We will need another important idea, first observed by Alon, Radoičić, Sudakov, and Vondrák [1] and Axenovich and Martin [2]: it is possible to find a $4$-colouring of $E(Q_{n})$ without monochromatic chordless $C_{10}$. More precisely, denote by $C_{6}^{-}$ any subgraph of $Q_{n}$ obtained by removing an edge from a copy of $C_{6}$ in $Q_{n}$. Note that every $C_{6}^{-}$ is a path with $5$ edges, but not every such path is a $C_{6}^{-}$. Define $\textup{ex}^{*}(Q_{n},C_{6}^{-})$ as the maximum number of edges in a subgraph of $Q_{n}$ containing no $C_{6}^{-}$. We make use of the following result of Axenovich, Martin, and Winter [3], which is based on the $3$-colouring given by Conder [10].
Theorem 1.2 (Lemma 16 in [3]).
$\textup{ex}(Q_{n},C_{10})\geqslant\frac{1}{3}\cdot\textup{ex}^{*}(Q_{n},C_{6}^{-}).$
Thus, the core of the proof is to show the following theorem.
Theorem 1.3.
For any $r,n\in\mathbb{N}$ with $r\leqslant n$, there exists a $C_{6}$-free induced subgraph $G_{r}$ of $L_{r}(n)$ with
$e(G_{r})>\frac{c}{2}\cdot e(L_{r}(n)).$
Proof of Theorem 1.1 from Theorem 1.3.
Take $G$ to be the union of the graphs $G_{r}$ given by Theorem 1.3 for every odd $r\leqslant n$, so that $G_{r}$ is disconnected from $G_{r^{\prime}}$ for every $r\neq r^{\prime}$.
Observe that every induced subgraph of a layer containing a $C_{6}^{-}$ also contains a $C_{6}$; thus, since $G_{r}$ is a $C_{6}$-free induced subgraph of $L_{r}(n)$, it is also free of $C_{6}^{-}$. This implies
$\textup{ex}^{*}(Q_{n},C_{6}^{-})\geqslant e(G)>\frac{c}{2}\sum_{r=1}^{\left\lceil n/2\right\rceil}e\big{(}L_{2r-1}(n)\big{)}=\frac{c}{4}\cdot e(Q_{n}),$
which, by Theorem 1.2, concludes the proof. ∎
2. Proof of Theorem 1.3
Similarly to the argument used in [12], we consider the vector space $\mathbb{F}_{2}^{r}$. Also we fix a nonzero vector $v_{0}\in\mathbb{F}_{2}^{r}$. For each $i\in[n]$, pick a vector $v_{i}\in\mathbb{F}_{2}^{r}\setminus\{0\}$ uniformly at random. For any subset $S$ of $[n]$ define the multiset of vectors $M(S)=\{v_{i}:i\in S\}$.
Recall that each vertex of our hypercube layer $L_{r}(n)$ is identified with a certain subset of $[n]$ of size $r$ or $r-1$. Define the sets
$B_{r}=\bigg{\{}S\in{[n]\choose r}:M(S)\text{ forms a basis of }\mathbb{F}_{2}^{r}\bigg{\}},$
$B_{r-1}=\bigg{\{}S\in{[n]\choose{r-1}}:\{v_{0}\}\cup M(S)\text{ forms a basis of }\mathbb{F}_{2}^{r}\bigg{\}},$
and denote by $G_{r}$ the induced subgraph of the layer $L_{r}(n)$ on $B_{r-1}\cup B_{r}$.
We will show that this (random) graph $G_{r}$ is $C_{6}$-free (deterministically) and has a large expected number of edges, and therefore has the desired properties with positive probability.
Claim 2.1.
$G_{r}$ is $C_{6}$-free.
Proof.
Suppose there is a copy of $C_{6}$ in $G_{r}$. There are exactly three different coordinates flipped by its edges, so it must form the middle layer of some $3$-dimensional subcube of $Q_{n}$. More precisely, by permuting the coordinates, we may assume that this copy of $C_{6}$ is exactly of the form $L_{r}(n)\big{[}A_{1}\cup A_{2}\big{]}$, where
$A_{1}=\big{\{}\{i\}\cup I:1\leqslant i\leqslant 3\big{\}}\qquad\text{and}\qquad A_{2}=\big{\{}\{i,j\}\cup I:1\leqslant i<j\leqslant 3\big{\}}$
for some $I\subset[n]\setminus\{1,2,3\}$ with $|I|=r-2$.
Observe that, since $\{1,2\}\cup I\in A_{2}\subset B_{r}$, the collection of vectors $\{v_{1},v_{2}\}\cup M(I)$ forms a basis of $\mathbb{F}_{2}^{r}$, so, in particular, the vectors in $M(I)$ are linearly independent. Therefore, after taking the quotient of $\mathbb{F}_{2}^{r}$ by $\operatorname{Span}(M(I))$ we obtain a vector space $V$ isomorphic to $\mathbb{F}_{2}^{2}$.
Define $x_{0},x_{1},x_{2},x_{3}\in V$ to be the images of $v_{0},v_{1},v_{2},v_{3}$, respectively, under the quotient map. Since $A_{2}\subset B_{r}$ we have that
$\{v_{i},v_{j}\}\cup M(I)\text{ forms a basis of $\mathbb{F}_{2}^{r}$ for any $1\leqslant i<j\leqslant 3$,}$
i.e.,
$\{x_{i},x_{j}\}\text{ forms a basis of $V$ for any $1\leqslant i<j\leqslant 3$.}$
Similarly, from $A_{1}\subset B_{r-1}$, by the same argument as above, we obtain
$\{x_{0},x_{i}\}\text{ forms a basis of $V$ for any $1\leqslant i\leqslant 3$.}$
In particular, this implies that $x_{i}\neq 0$ for each $i\in\{0,1,2,3\}$ and that $x_{i}\neq x_{j}$ for any $i\neq j$. But $|V\setminus\{0\}|=3$, which yields a contradiction. ∎
Claim 2.2.
$\mathbb{E}\big{[}e(G_{r})\big{]}>\dfrac{c}{2}\cdot e\big{(}L_{r}(n)\big{)}$.
Proof.
Consider an edge of $L_{r}(n)$ connecting two sets $x=\{j_{1},\ldots,j_{r-1}\}$ and $y=\{j_{1},\ldots,j_{r}\}$. Define the vector spaces
$V_{0}=\operatorname{Span}\{v_{0}\},\;V_{k}=\operatorname{Span}\{v_{0},v_{j_{1}},\ldots,v_{j_{k}}\}\;\text{ for $1\leqslant k\leqslant r-1$,}$
[…]
Note that
$\mathbb{P}\big{(}xy\in G_{r}\big{)}=\mathbb{P}\big{(}x\in B_{r-1}\text{ and }y\in B_{r}\big{)}=\mathbb{P}\big{(}\dim V_{r-1}=\dim V_{r}=r\big{)}.$ (1)
Observe that, if the vectors $v_{0},v_{j_{1}},\ldots,v_{j_{k-1}}$ (for $11\leqslant k\leqslant r-$ 1) are fixed and linearly independent, then there are exactly $2^{r}-2^{k}$ choices for the next vector $v_{j_{k}}$ such that $v_{0},v_{j_{1}},\ldots,v_{j_{k}}$ are still linearly independent, as it is equivalent to the condition $v_{j_{k}}\notin V_{k-1}$. Similarly, if the vectors $v_{j_{1}},\ldots,v_{j_{r-1}}$ are fixed and linearly independent, then there are $2^{r}-2^{r-1}$ choices for the last vector $v_{j_{r}}$ such that $v_{j_{1}},\ldots,v_{j_{r}}$ are still linearly independent, as it is equivalent to the condition $v_{j_{r}}\notin\operatorname{Span}\{v_{j_{1}},\ldots,v_{j_{r-1}}\}$. Therefore, by (1), we have
$\mathbb{P}\big{(}xy\in G_{r}\big{)}$ $=\Bigg{(}\prod_{k=1}^{r-1}\mathbb{P}\big{(}v_{j_{k}}\notin V_{k-1}\;|\;\dim V_{k-1}=k\big{)}\Bigg{)}\cdot\mathbb{P}\big{(}v_{j_{r}}\notin\operatorname{Span}\{v_{j_{1}},\ldots,v_{j_{r-1}}\}\;|\;\dim V_{r-1}=r\big{)}$
$=\Bigg{(}\prod_{k=1}^{r-1}\frac{2^{r}-2^{k}}{2^{r}-1}\Bigg{)}\cdot\frac{2^{r}-2^{r-1}}{2^{r}-1}>\frac{1}{2}\prod_{k=1}^{\infty}\bigg{(}1-\frac{1}{2^{k}}\bigg{)}=\frac{c}{2}.$
The claim now follows from linearity of expectation. ∎
As observed above, it follows from Claims 2.1 and 2.2 that there exists a choice of the vectors $v_{1},\ldots,v_{n}$ such that
$e(G_{r})>\frac{c}{2}\cdot e\big{(}L_{r}(n)\big{)}$
and $G_{r}$ is $C_{6}$-free, as required.                                                                                                                                     ∎
Acknowledgements
We would like to thank the anonymous referees for their valuable suggestions. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior, Brasil (CAPES).